OpenStudy (anonymous):

Let z be a complex number such that the imaginary part of z is nonzero and A = z^2 + z + 1 is real. Then A cannot take the value of what?

OpenStudy (zarkon):

What have you tried?

OpenStudy (anonymous):

i wrote z as a+bi and then plugged that into z^2+z+1 @Zarkon

OpenStudy (anonymous):

i got \[a^2+2abi-b^2\]

OpenStudy (anonymous):

so how do i find it what it can;t take the value of?

OpenStudy (anonymous):

if a^2+2abi-b^2 is real?

OpenStudy (zarkon):

you get more than that...you didn't include the z+1 part

OpenStudy (anonymous):

oh yah oopps

OpenStudy (anonymous):

so its \[a^2+2abi-b^2+a+bi+1\]

OpenStudy (anonymous):

which if i simplify is\[a^2+(2a+1)bi-b^2+a+1\]

OpenStudy (anonymous):

i factored out a 2a+1 because 2a+1 must be real right?

OpenStudy (anonymous):

so that means that (2a+1)bi is complex right?

OpenStudy (zarkon):

if A is real then (2a+1)bi=0

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

oh yah there is no imaginary part right?

OpenStudy (anonymous):

oh so it simplifies to:\[a^2-b^2+a+1\]

OpenStudy (anonymous):

now what should i do?

OpenStudy (anonymous):

\[(a+b)(a-b)+a+1\] this?

OpenStudy (anonymous):

ok i think i can do it from here @Zarkon

OpenStudy (anonymous):

no actually i'm stuck :(

OpenStudy (anonymous):

i also wrote it like:\[a(a+1)-b^2+1\]

OpenStudy (anonymous):

well -b^2 will always be negative

OpenStudy (zarkon):

if (2a+1)bi=0 and b is non zero then a=-1/2

OpenStudy (anonymous):

thanks @Zarkon